### v1

# ALIGNING AUDIO CAPTIONS WITH HUMAN PREFERENCES

Abstract
Current audio captioning systems rely heavily on supervised learning with paired audio-caption datasets, which are expensive to curate and may not reflect human preferences in real-world scenarios. To address this limitation, we propose a preference-aligned audio captioning framework based on Reinforcement Learning from Human Feedback (RLHF). To effectively capture nuanced human preferences, we train a Contrastive Language-Audio Pretraining (CLAP)-based reward model using human-labeled pairwise preference data. This reward model is integrated into a reinforcement learning framework to fine-tune any baseline captioning system without relying on ground-truth caption annotations. Extensive human evaluations across multiple datasets show that our method produces captions preferred over those from baseline models, particularly in cases where the baseline models fail to provide correct and natural captions. Furthermore, our framework achieves performance comparable to supervised approaches with ground-truth data, demonstrating its effectiveness in aligning audio captioning with human preferences and its scalability in real-world scenarios.
Index Terms—  Audio Captioning, preference alignment, reward, RLHF
1 Introduction
Audio captioning, the task of generating natural language descriptions for acoustic scenes, has become an important benchmark for evaluating audio-language models. Unlike image captioning, which benefits from explicit spatial and semantic cues, audio captioning faces inherent ambiguity and temporal complexity, making accurate and human-aligned caption generation particularly challenging.
While supervised approaches using datasets such as AudioCaps [1] and Clotho [2] have advanced the field, their reliance on automatic metrics like BLEU [3], METEOR [4], and CIDEr [5] introduces limitations, as these metrics correlate poorly with human judgments [6]. Moreover, recent contrastive learning methods such as CLAP, though scalable, fail to capture nuanced human preferences, underscoring the need for approaches that optimize for human alignment rather than static similarity metrics.
Optimizing task-specific metrics using Reinforcement Learning (RL) was popularized in image captioning by Self-Critical Sequence Training (SCST) [7]. However, metric-based RL optimization remains underexplored in the domain of audio captioning. Recent work [8] has demonstrated improved alignment with human expectations by leveraging a CLAP-based reward model combined with a large language model, outperforming traditional CIDEr-based RL approaches on mean-opinion score based human evaluation. Despite these gains, the reliance on external language models introduces significant computational overhead. Also, evaluation of audio captions with absolute rating can cause poor inter-annotator agreement [6].
To address the challenges associated with audio captioning such as the reliance on paired audio-caption ground truth data and the limitations of existing evaluation metrics we propose a framework that enhances caption quality based on human preferences. Our method does not require additional paired data, making it more scalable and cost-effective. Since collecting preference data is significantly easier and less resource-intensive than obtaining ground-truth captions, we design a simple yet effective reward model that leverages pairwise human judgments. To ensure training stability and mitigate reward hacking during reinforcement learning, we incorporate reward shaping techniques [9]. We conduct a comparative study between captions generated by the baseline model and those produced by our proposed method, evaluated both through human assessments and established automatic metrics. Additionally, we analyze the performance of our approach relative to supervised training methods.
Fig. 1: Pairwise human preference annotation and subjective evaluation setup
Fig. 2: Proposed Reward Model architecture
Fig. 3: Proposed RL Pipeline
Our key contributions are:
  • We propose a novel audio captioning framework that optimizes a custom reward function using reinforcement learning, without requiring paired audio-caption data.
  • We develop a CLAP embedding-based reward model trained on pairwise human preference data to effectively capture subjective judgments.
  • We conduct extensive human evaluations on publicly available and proprietary datasets, showing that our RLHF-based fine-tuning consistently outperforms baseline models, especially when the baseline model does generated captions are incorrect or unnatural. Additionally, our method achieves performance comparable to supervised training, while requiring no labelled audio-caption data.
2 Related Work
Early works in audio captioning, such as AudioCaps [1] and Clotho [2], introduced large-scale datasets pairing audio clips with human-written captions, enabling supervised fine-tuning (SFT) of captioning models. Evaluation has traditionally relied on image captioning metrics like BLEU [3], METEOR [4], and CIDEr [5], but these metrics poorly capture semantic fidelity and subjective quality [6, 10]. Alternatives such as FENSE [6, 11] combine semantic similarity with error modeling for improved human alignment.
CLAP-based models [12, 13] learn joint embeddings for audio and text, and CLAPScore has been used as a proxy for caption relevance and as a hallucination metric [14]. However, recent studies indicate that CLAPScore correlates only weakly with human judgments [15], and that CLAP embeddings suffer from a modality gap that limits fine-grained semantic alignment [16]. These findings suggest that optimizing solely for CLAP similarity may produce captions that are technically relevant but lack naturalness and contextual depth. To address this, we adapt a neural network over CLAP embeddings to better capture human preferences.
RLHF has emerged as a powerful paradigm for aligning model outputs with human expectations, particularly in large language models. Self-Critical Sequence Training (SCST) [7] applies policy gradient methods to optimize non-differentiable metrics like CIDEr. Alternative approaches such as Direct Preference Optimization (DPO) [17] have shown promise in text domains but, due to reliance on static pairwise comparisons and limited control over output structure [18], they are less suited for concise audio captioning systems. Our work builds on SCST and extends RLHF to audio captioning with a custom reward model and reward shaping strategies to stabalize training [8, 9].
3 Methodology
In this section we describe about reward modelling for custom preference data and fine-tuning Audio Captioning model with reinforcement learning. An overview of our proposed approach is shown in Figure 3.
3.1 Reward Modelling
The goal of the reward model is to capture the subtleties in human preference data by assigning a scalar value that estimates the probability of a human preferring a given audio-caption pair. Unlike reward models for language tasks which typically use a pretrained language model to generate multiple completions for a given prompt, our approach requires a multimodal reward model. To achieve this, we utilize the CLAP audio and text encoders as described in [13] to extract 512-dimensional embeddings for both modalities. These embeddings are concatenated to form a joint 1024-dimensional representation, which is then input into a simple two-layer neural network with hidden layer sizes of 512 and 128, respectively. The final output layer has a single neuron mapped to the range [0, 1] using sigmoid activation.
The preference dataset comprises tuples of the form $(x,y_{1},y_{2},p)$ where $x$ is the audio input, $y1$ and $y2$ are two candidate captions, and $p$ indicates the human-annotated preference. As per Bradley-Terry model [19], a widely used approach for modeling pairwise preferences, the probability that caption $y1$ is preferred over $y2$ for a given audio input $x$ is given by:
$P(y_{1}\succ y_{2}\mid x)$ $=\frac{e^{r(x,y_{1})}}{e^{r(x,y_{1})}+e^{r(x,y_{2})}}$
$=\sigma(r(x,y_{1})-r(x,y_{2}))$ (1)
where $r(x,y)$ denotes the reward assigned by the model to the audio-caption pair, and $\sigma$ indicates the sigmoid function.
The reward model has a siamese network architecture. For the preferred caption $y_{w}$ and the non-preferred caption $y_{l}$, the model computes rewards $r_{\theta}(x,y_{w})$ and $r_{\theta}(x,y_{l})$ respectively. We scale the difference of reward by a factor of $\beta$. To prevent overfitting and encourage stable reward predictions, we incorporate an L2 regularization term into the loss function with the regularization coefficient $\lambda$, resulting in the following loss function:
$\mathcal{L}(\theta)=\mathcal{L}_{\text{BT}}(\theta)+\mathcal{L}_{\text{reg}}(\theta)$ (2)
$\mathcal{L}_{\text{BT}}(\theta)=-\log\left[\sigma\left(\beta\left(r_{\theta}(x,y_{w})-r_{\theta}(x,y_{l})\right)\right)\right]$ (3)
$\mathcal{L}_{\text{reg}}(\theta)=\lambda\cdot\left[r_{\theta}(x,y_{w})^{2}+r_{\theta}(x,y_{l})^{2}\right]$ (4)
During inference, as shown in Figure 3, the reward model receives an audio-caption pair as input and outputs a score between 0 and 1, which serves as a proxy for the probability of human preference.
Table 1: Metric-wise comparison of baseline and proposed RLHF systems
Dataset Test Data System N-gram Semantic Preference based Proposed RLHF Win rate (%)
BLEU4 CIDEr ↑ S-BERT↑ FENSE ↑ CLAPAT Reward ↑ Reward ↑ Human ↑
AudioCaps Non-challenging Baseline 0.0272 0.3211 0.5947 0.5947 0.3517 0.6524 48.20 52.03
+RLHF 0.0314 0.3178 0.5836 0.5282 0.3231 0.6480 51.80 47.97
Challenging Baseline 0.0541 0.3357 0.4565 0.1441 0.2546 0.5482 40.86 46.07
+RLHF 0.0498 0.3465 0.5007 0.3571 0.2572 0.6025 59.14 53.93
Proprietary Non-challenging Baseline 0.0919 0.7970 0.5392 0.4772 0.3015 0.1958 35.49 46.59
+RLHF 0.0771 0.6324 0.5371 0.4688 0.2820 0.2981 64.51 53.41
Challenging Baseline 0.0780 0.4570 0.5194 0.4637 0.2425 0.1558 30.12 40.89
+RLHF 0.0555 0.4463 0.5607 0.4933 0.2282 0.3128 69.88 59.11
3.2 Reinforcement Learning
Our objective is to enhance the baseline audio captioning model to generate captions that are more closely aligned with human preferences. We formulate this as a reward optimization problem, where the reward predicted by our reward model is maximized to obtain improved captions. Since the scalar reward is non-differentiable and direct gradient computation and model updates are not feasible, we employ the Policy Gradient method using REINFORCE [20]. To further reduce the variance of the gradient estimate, a baseline term is introduced, which does not affect the expected gradient. We employ Self-Critical Sequence Training [7] approach, which uses the output of the test-time inference algorithm specifically, greedy decoding, as the baseline.
To compute the loss, we consider reward scores for two captions generated by the baseline model using different decoding methods. Using our custom reward calculator not only provides more control than using the traditional metrics such as CIDEr, it also eliminates the need for ground truth captions to compute the reward score, hence affording scalability. The captions $w^{s}$ (decoded using multinomial sampling) and $w^{g}$ (decoded using greedy decoding) along with the audio are passed to the reward model to obtain rewards, denoted as $r(w^{s})$ and $r(w^{g})$. These reward values are then used to compute the loss and estimate the gradient, given by:
$\nabla_{\theta}L(\theta)\approx-(r(w^{s})-r(w^{g}))\nabla_{\theta}\log p_{\theta}(w^{s})$ (5)
A known challenge in reinforcement learning (RL) for text generation is reward hacking, where models learn to exploit the reward function in ways such as generating excessively long captions or sometimes repiting certain words that superficially maximize the reward without improving the semantic quality or relevance.
3.2.1 Length Penalty
During our initial experiments, we observed that length based reward hacking was dominant. To address reward hacking, we apply reward shaping by introducing a length penalty aligning with recent work by [9], who demonstrate how reward shaping can mitigate reward hacking.
We define the modified reward function as:
$r_{\text{new}}=r_{\text{old}}-\alpha\cdot\left(1-\frac{L_{e}}{L_{c}}\right)\cdot\max(0,L_{c}-L_{e})$ (6)
Here, $L_{e}$ is the expected caption length, $L_{c}$ is the actual generated caption length, and $\alpha$ is a tunable hyperparameter controlling the strength of the penalty. The penalty gets applied only when the generated caption length exceeds the expected length of caption. This formulation when used in equation 5, penalizes overly long captions proportionally to their deviation from the expected length, helping to align generation with user preferences.
4 Experimental Setup
We conduct our experiments on both publicly available and proprietary datasets to evaluate the performance of our proposed system.
Table 2: Qualitative comparison of Baseline, RLHF, SFT, and human labelled Ground Truth captions on AudioCaps dataset
Baseline RLHF SFT Ground Truth
A musical genre is playing A musical instrument is playing Music plays - A piano playing as plastic bonks
- A piano playing as a clock ticks
- A clock ticking as a piano note is playing
A man is talking while a person is snoring A man is talking while there is a continuous buzzing of an insect A man speaks and a woman laughs - Bees buzz while two men talk followed by a sniffle
- Swarm of bees buzzing and two men speaking
- Men speaking with insects buzzing
4.1 Dataset
To train the reward model, we construct a preference dataset comprising audio samples, caption pairs, and human-annotated preferences based on caption correctness and naturalness.
Due to limited availability of public audio-caption preference datasets, we utilize AudioCapsEval and ClothoEval from FENSE [6], derived from AudioCaps [1] and Clotho [2]. These datasets contain approximately 1,750 annotations each, with caption pairs rated by four human annotators. We extract 1,473 pairs from AudioCapsEval and 1,555 from ClothoEval, retaining only samples with unanimous annotator agreement. The data is split 80% for training and 20% for validation.
To improve model performance on proprietary systems, we curate a custom dataset of 4,424 preference samples and 880 challenging cases. For each audio sample, captions are generated via greedy decoding and top-$k$ sampling. Annotators select the preferred caption or provide a corrected one if both are incorrect. Samples are labeled challenging when both captions fail. As public datasets lack such annotations, we approximate challenging cases using FENSE scores [6], selecting the bottom 250 samples from AudioCapsEval.
To enhance diversity and simulate misaligned scenarios, we augment the proprietary dataset with synthetic preference pairs. Each audio sample is paired with a randomly selected caption from a different sample, creating semantically irrelevant ‘losing’ candidates. This helps the reward model learn to penalize mismatched captions.
4.2 Training and implementation details
We consider the model proposed in [21] as our baseline for all experiments which employs a CNN10 PANN [22] as the audio encoder and a transformer decoder. The model is trained in a supervised manner on 80,000 proprietary audio-caption pairs and comprises approximately 15 million parameters.
We use pre-trained LAION clap-htsat-unfused model [13] for extracting audio and text embeddings as input features. The reward model was trained for 70 epochs with a batch size of 64 and a learning rate of $10^{-4}$. To stabilize training and avoid extreme reward values, output predictions were clamped within the range [0.01, 0.99]. We scale the difference in the reward by a factor $\beta$ equals to 5. L2 regularization was applied via a coefficient of 0.1 to prevent overfitting.
During RLHF, length penalties of 0.4 and 1.0 were applied for the AudioCaps and proprietary datasets respectively, with an expected caption length $L_{e}$ of 13. We perform RLHF for 100 epochs, with a batch size of 128, using the Adam optimizer with a learning rate of $10^{-6}$ and an L2 weight decay of $10^{-6}$. A warmup phase linearly increases the learning rate over the first two epochs with a multiplier of 1.1, after which, a step decay schedule reduces the learning rate by a factor of 10 every 10 epochs.
Table 3: Comparison of Reward Models
Reward Model Data Training Data Size Human Win Rate for RLHF
AudioCapsEval 1178 43.55%
AudioCapsEval + ClothoEval 2422 47.97%
4.3 Human Evaluation
To evaluate the models, we collected human feedback between pairs of audio captions in a manner similar to [6]. We asked annotators to listen to the audio and provide preference based on 2 given captions. For unbiased evaluation, we randomized the order of caption generated by two different methods. To compare RLHF captions with baseline captions, we had four annotators indicate their preferences for 250 caption pairs. For the ablation study, to compare RLHF captions with SFT captions, we had two annotators to indicate their preferences for 100 caption pairs. Human evaluation can vary from person to person, to validate if there are inter-annotator agreement between the users, we calculate the Fleiss Kappa score [23]. Based on our subjective evaluation, it is found that for AudioCaps dataset, agreement is 0.445 and 0.435 over non-challenging and challenging data and for Proprietary dataset, agreement is 0.406 and 0.417 over non-challenging and challenging data, which is very acceptable. We consider a preference to be win when caption generated by proposed model is preferred over oher caption, and the win rate is the percentage of preferred RLHF over total non-zero preference.
5 Results and Discussion
We evaluate our RLHF framework on both public and proprietary datasets, including challenging cases where baseline models fail to produce accurate or natural captions. Captions are assessed using n-gram-based metrics, semantic-aware metrics, our custom reward model, and human preference judgments. Table 1 shows that even though the baseline model performs well on automatic metrics, it fails to align with human preferences highlighting the limitations of rule-based and semantic metrics. In contrast, our RLHF model consistently improves human preference win-rates, except on the AudioCaps evaluation set, likely due to limited preference data for training the reward model. Preference based reward model correlates well with human judgments, and RLHF maintains competitive scores on automatic metrics. Despite being built upon the $CLAP$ model, our reward model exhibits a more consistent similarity trend with human evaluations than the conventional $CLAP_{Audio\text{-}Text}$ score. Human preference win-rate analysis reveals that when the baseline model fails to generate high-quality captions, the reward model assigns lower scores to those samples, thereby steering the learning process toward outputs that better reflect human judgment. Table 2 shows two scenarios where proposed RLHF aligns well with humans either based on naturalness or the correctness of the caption, compared to the other methods.
5.1 Ablation study
Table 3 illustrates that increasing the amount of preference data leads to a more effective reward model, which in turn enhances RL fine-tuning. This improvement results in higher-quality caption generation that aligns more closely with human preferences.
Fig. 4: SFT v/s RLHF comparison
Figure 4 compares different training methodologies. For the AudioCaps dataset, supervised fine-tuning was performed using additional ground-truth captions from the train split, while proposed RLHF only uses audio from the same AudioCaps training data. In contrast, for the proprietary dataset, we used the winning captions from the preference dataset as ground-truth for supervised tuning and only audio for the RLHF fine-tuning. Due to limited human annotaotrs, we conducted human study with 2 annotators asked to annotate 100 samples for each experiment. The results indicate that supervised fine-tuning benefits from the availability of paired ground-truth data, leading to improved performance. However, the proposed RLHF approach remains competitive, achieving comparable results despite being fine-tuned without access to additional ground-truth annotations.
6 Conclusions
We presented a reinforcement learning framework for audio captioning that aligns model outputs with human preferences using a custom reward model trained on preference data. Our RLHF-based approach improves human preference alignment without relying on ground-truth captions, while maintaining competitive performance on traditional metrics. The method is scalable, simple to implement, and effective in challenging scenarios, with further gains possible through enhanced reward modeling with expanded preference data and advanced RL techniques.